FEL Upgrade Project Weekly Brief
October 15-19, 2001






Highlights:
Results from last week's measurements of the length of the FEL pulse show that the optical pulse length can be shortened to as low as 180 fs (FWHM).

The first evidence for FEL excitation of a molecule in a molecular beam was obtained by scientists from the UVA Chemistry Department.  Future experiments will provide valuable insight into chemical dynamics.

The drawings and manufacturing specifications for the largest magnets in the Upgrade, the 180° recirculation dipoles (DY),were signed and the package was sent to Procurement.

Management:
The agenda for the Nov. 8-9th semiannual review of the FEL Upgrade project was reviewed with the ONR contract monitor and distributed to the review team and prospective attendees.   Members of FEL team continued their preparations for presenting summaries of the technical, cost and schedule status of each Upgrade subsystem at the review.

The JLab and Aerospace procurement organizations successfully completed negotiations for our major subcontract under the first phase of the UV FEL project.  We expect to send the package to DOE next week for approval.

Bill Nay, who oversees security issues for DOE's Office of Science, visited the Lab on Monday (Oct. 15th) to discuss potential security/export control implications of the lab's activities in the design and operation of high average power free electron lasers.

WBS 3 (Beam Physics):
DIMAD was ported to WinNT and is being tested.  BBU threshold simulations are being set up for the Revision 1.1.2 design. Documentation of the Revision 1.1.2 design is nearly complete.

WBS 4 (Injector):
The ceramic which had a leak in the braze joint was cut off the stack and preparations for welding a new one on are being done.  The large tube for the main gun chamber has been machined and is being sent out to have the I.D. electro-polished.  A test section of the wire for feeding power to the cesium channels inside the ball cathode was made and works very well.  The drawings are being fine tuned prior to sign off.  Concepts of a new resistor design allowing for the removal of the elbow from the high voltage transmission line (would increase reliability) are being looked at.

Gun HVPS - Plans are starting to move the HVPS tank to be able to eliminate the elbow and the transmission line hangers.  T. Siggins is looking at making a new assembly that combines the running and conditioning resistors.

WBS 5 (SRF):
Received second space frame and vacuum vessel from the vendor.  These include fixes and modifications as identified in the first article.  Inspection and acceptance is in process.

WBS 6 (RF):
Zone 3 - Completed with only a few items on a punch list.

Zone 4 - Completed except for position 8 needing its circulator and klystron tested and calibrated.

Injector RF - Ion pump controllers were ordered this week for the 100 kW klystrons.  The lifting plate is designed and parts are on order.  The designed will be approved before the part is built.

Injector HVPS - The design of the HVPS is nearly complete with the specification of the HV inductor and capacitors.  Quotes are coming in for the purchase of the capacitors.  The layout of the HVPS major components is in progress.

WBS 8 (Instrumentation):
Final parts for the beamviewer camera/solenoid power supply boxes have been received and assembly has began.

A bar coding program is being developed for component tracking during the rip-out and subsequent re-installation.  This will also translate into an inventory control system and cable tracking system as the new machine is installed.  Users will be able to request new labels by filling out a form on the web.  The labels will then be printed by a dedicated printer for pick-up and attaching to the component.

Studies were done cold cathode vacuum controllers for both the upgrade and the SNS project. These gauges will be the vacuum interlock for the input couplers. The absolute pressure and time response to trips were measured, these can be seen at:  http://laser.jlab.org/sns.

WBS 9 (Transport):
Dipoles
Optical Chicane Dipoles (GW)
o  Bulk steel, copper and pole tip sheet steel were received by Magnet Enterprises International.
o  We  plan to inspect the first coil potting in four weeks.
Injector Dipoles (DU/DV)
o  The drawings went through final review.  Some small portions of the final assemblies need modification.  The
    details are fine.
Arc Dipoles (GY, GX, GQ)
o  Design continued on backchecking the detail drawings of the bend and Reverse Bend Dipoles (GX & GQ)
    with the comments and corrections.
o  The Drawings and manufacturing specifications for the 180° Dipole (DY) were signed and the package was
    sent to Procurement.
o  We  notified a number of manufacturers throughout the country about the coming procurement and made
    preliminary drawings and specifications available to them .
Magnet Measurement
o  The drawings for the measurement stands for the two quadrupoles were signed and released for manufacture.
o  The design of the dipole measurement stand continued with the all important probe drawing nearing
    completion.  A final resolution of the measurement of the GX dipole's UV branch held up final approval.
o  We started work on the probe that will measure the sextupole and the octupole.
Quadrupoles
3 inch quad (QX)
o  New England Technicoil, the core and assembly vendor had the first article core inspected by Al Guerra this
    coming Monday.  (Al was sick and his inspection reported last week was delayed one week) The first batch
    of rough machined cores had gone through annealing. New England Technicoil moved  up their schedule for
    shipment of the first batch from December  to November.
Trim Quad (QT)
o  The bid for this magnet has gone back out for a quick bid cycle to make available Jefferson Lab supplied
    steel.  Some vendors were having difficulty getting that grade.
Sextupole (SF)
o  Work  on the magnetic model was interrupted because of cooling problems in the septum magnets of the
    CEBAF Accelerator.
Octupole (OT)
o  DULY Research completed their preliminary mechanical model of the octupole with an excellent analysis of
    the required shape of the pole tip and field clamp to achieve David Douglasí specifications.  They have now
    been diverted to Light Box design.
Corrector Dipoles (DB, DJ)
o  Milhous Co. shipped the first batch of coils.  They look good.  They are under notification to complete the
    contract by Thanksgiving.
Beam Line and Vacuum
o  Dave Waldman generated the preliminary task order for design of the Arc Chambers utilizing the Engineering
    Services contract with Advanced Energy Systems and they have given us feed back.
o  The bids for girders and pedestals necessary for the return leg and the hardware for the entire machine were
    remain in evaluation.
o  Design work on girders intensified with a major layout of the quadrupole telescope in front of the first arc and
    the region between the cryomodules  heading towards detailing.
o  The modifications of the injector line stands and girder for dipoles have been finalized.

WBS 10 (Wiggler):
Production runs have started for wiggler field measurements.  We were able to reduce the level of noise in the system substantially by adding a grounded sheath to the Hall probe cable.  The vacuum chambers are complete and leak checked with the exception of the two flanges which mate to the dispersion section.  These will be welded when fitup is accomplished.  The aperture of the chamber was checked and it was found adequate for operation out to 25 microns if a Rayleigh range of 340 cm is used and the cavity waist is in the center to within 30 cm.

WBS 11 (Optics):
We received the rest of the fabricated parts for the ultraviewers.  We'll begin assembly and testing next week.  Detailing continues on the optical cavity.  Particular attention was devoted to the sizing of the vertical translation rails.  After the internal review, we though they needed to be resized, but this would have considerably changed component placement.  With some changes in rail separation and component stack up, the design team came up with a solution that drops the rail diameters to the original size(due to reduced torque).   They have also slightly redesigned the yaw mount to make it stiffer.  We learned that the lab has had some problems with using Dicronite (tungsten disulfide) as a high vacuum-compatible low-friction coating, and so we're looking at other options.

The following work was done to support operations:

We provided considerable assistance to the UVa team with installation of optics and position diagnostics for their experiment.  We provided optics and an IR camera for the NSU experiment, and a linear stage and hardware for the X-ray experiment.  Based on our experience with the measurements made in the Mirror test Stand (MTS) last week, and the availability of parts, we changed the experimental configuration slightly to get better data.  We replaced the zoom motor on the Wyko 4100 laser interferometer, and recalibrated the zoom encoder.  We have additional near-Brewster windows ready for installation in User Labs 2 and 4 at the first opportunity.  Transitions from these new windows to the experiment hutches are being fabricated.

Operations/Commissioning:
We lost some experimental time this second week of the Fall user run due to a break in the low conductivity main feed on the accelerator site.  The FEL building was isolated from the break and we resumed operation on late Wednesday evening.

The first attempts by members of the UVa Chemistry Dept. to look at vibrational excitation of molecules in a molecular beam were executed this week.  Using bolometric detection of molecules of pyrone in a supersonic nozzle beam, a strong signal was observed.

We are currently focusing on signal to noise issues.  A Michelson arrangement will later be used to generate 2 beams with a variable time delay for intramolecular dynamics experiments.

We continued to make substantial quantities of carbon nanotubes and to study production dependencies.

An additional effort at near field microscopy at 3 microns was attempted but was unsuccessful.

Work on Thompson X-ray studies continued as a background study.